# Black Hole Survival Guide: Searching for Stars in the Galactic Center That Endure Partial Tidal Disruption

### v1

@@ 3.2 Colors and Abundances
The altered composition of the remnant is especially noticeable for the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant, shown in the right-most panel of Figure 5. Compared to a solar-mass star at solar abundance, which has a similar total mass to this remnant, each of the hydrogen, helium, carbon, and nitrogen abundances at the surface are found to be appreciably altered ($X_{\rm He}/X_{\mathrm{He},\odot}\approx 2$, $X_{\rm N}/X_{\mathrm{N},\odot}\approx 7$, $X_{\rm C}/X_{\mathrm{C},\odot}\approx 0.2$). This is still caused by convective mixing after the TDE, but the effect is enhanced since the $3\,M_{\odot}$ progenitor primarily burned hydrogen on the main sequence via the CNO cycle before the TDE occurred. For these highly disrupted, higher-mass main sequence stars, we expect an even greater prospect of distinguishing a history of tidal disruption through an unusual stellar spectrum. Our findings are corroborated by a similar study completed by M. Sharma et al. (2024), in which they simulate the evolution of TDE remnants of $M_{i}=1$ and $M_{i}=3\,M_{\odot}$ MAMS stars and also find that high-$\beta$ remnants exhibit compositional mixing that leads to heightened surface enrichment of N and He (Figures 13, 14, and 18 in M. Sharma et al. 2024).
3.3 Long-Term Stellar Evolution
Our three remnants disrupted from a $1\,M_{\odot}$ progenitor lose varying degrees of mass during the tidal disruption event and consequently spend the rest of their lives burning less intensely. This lowered rate of fusion resulting from their mass loss allows them to live billions of years longer than their progenitor. While the $1\,M_{\odot}$ progenitor ages off of the main sequence around age 11 Gyr (roughly 6 Gyr after disruption), the $\beta=1$, $M_{i}=1\,M_{\odot}$ remnant’s life is extended by an additional 3 Gyr; by 19 Gyr for the $\beta=1.5$, $M_{i}=1\,M_{\odot}$ remnant; and by 140 Gyr for the $\beta=2$, $M_{i}=1\,M_{\odot}$ remnant, as seen in Figure 2. Conversely, each remnant has a shorter lifetime than its mass-equivalent star, which is consistent with the remnant having the core density of a higher-mass star.
Once the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant reignites core hydrogen fusion, it similarly lives longer than its progenitor and reaches the RGB earlier than its mass-equivalent star. Notably, the remnant’s luminosity during hydrogen burning is larger than that of its mass-equivalent star by a factor of $\approx 10$ during its long-term evolution on the main sequence, which is a much larger contrast than any of the remnants disrupted from a $M_{i}=1\,M_{\odot}$ progenitor achieve. Consequently, it may be possible to differentiate the TDE remnant of a heavily disrupted high-mass star from the remnant of a mildly disrupted low-mass star, in that a more heavily disrupted high-mass remnant maintains a heightened luminosity signature for its entire life, even after having reignited core hydrogen fusion.
While our work is most concerned with the previously unstudied early-time behavior of the remnants, we evolve our models self-consistently until the end of their time on the main sequence to offer a wholistic picture of their entire evolution from initial conditions to old age. We find that our early-time behavior leads naturally to late-time behavior that is consistent with M. Sharma et al. (2024), which studied the late-time evolution in detail. At late times, both of our studies find that the remnant is cooler and dimmer than its progenitor, yet hotter and brighter than its mass equivalent star; that the remnants resulting from higher-$\beta$ interactions live longer; and that the brightest remnants come from low-$\beta$ encounters that lose the least amount of mass.
We note that our study differs from M. Sharma et al. (2024) in that our hydrodynamical FLASH calculations have 10$\times$ higher resolution. This allows us to effectively capture the early evolutionary behavior for $t\lesssim 10^{4}$ years, during which time the TDE remnant is brightest and may be most visible. There is value in modeling the early evolution of the remnant, as it provides a sense of how the mixing evolves over time, showcases their brightest and hottest phase, and opens the question of whether certain rare objects and bright transients might be explainable as young TDE remnants.
@@ 4 Discussion and Conclusions
In this work, we investigate the post-disruption evolution of partial TDE remnants in order to develop strategies for discovering these objects. Our models indicate the remnant will be brightest in the first $10^{5}$ years after disruption, during which period they may appear unusually luminous and extended compared to undisturbed stars. Over timescales of $\approx 10^{7}$$10^{8}$ yr, however, the majority of TDE remnants will have long faded to appear quite similar to main sequence stars of a similar mass. The (relatively uncommon) high-$\beta$ remnant from a more massive progenitor does appear substantially offset from its mass-equivalent, but all remnants are still expected to fall within the natural variation of MS stars on Hertzsprung-Russell diagram. Nevertheless, TDE remnants constitute a class of interesting objects which may be identified observationally in a few key ways.
@@
We are particularly interested in similarities between our models of TDE remnants and the known class of G objects in the Galactic Center. The G objects move with star-like kinematics, but have massive dust-shrouds that make them appear more like clouds. Though they are yet to be explained, the G objects pose interesting candidates for young TDE survivors, given their location in the Galactic Center, elevated luminosities, and puffy radii (Figure 2).
@@
In order to understand whether TDE remnants may be related to G objects, it is important to estimate whether TDEs occur at a rate high enough to suggest that young, high-luminosity remnants could account for some or all of the current population of G objects. Partial TDEs are expected to occur at larger rates than full disruptions (E. Ramirez-Ruiz & S. Rosswog, 2009; J. Guillochon & E. Ramirez-Ruiz, 2013; M. MacLeod et al., 2013); for instance, in the full loss cone regime, the rate of TDEs should scale with the inverse of $\beta$ (M. MacLeod et al., 2012; J. Krolik et al., 2020; N. C. Stone et al., 2020). The heightened partial TDE occurrence rate is countered by the fact that our models show partial TDE remnants fading rapidly within $\lesssim 10^{5}$ yr. Accounting for both effects, we estimate that at any given moment in the Galactic Center, we can expect a handful of TDE remnants to exist that have high enough luminosities to match those of the dimmest G objects. Thus, we predict that TDE remnants in their early phase of evolution could account for a few G objects, but not the entire population (B. N. Sitarski, 2016).
Given that the vast majority of TDE remnants have aged past their initial energized state, what are the best criteria to use when attempting to identify an older remnant after it has faded? We recommend employing a combination of kinematics and spectroscopy. TDE remnants on elliptical orbits remain bound to the Galactic Center (H. Manukian et al., 2013) and may experience repeated disruptions by the SMBH (M. MacLeod et al., 2013; M. Cufari et al., 2022; C. Liu et al., 2023; A. Bandopadhyay et al., 2024; C. Liu et al., 2025). However, those remnants that maintain parabolic or hyperbolic orbits will eventually exit the Galactic Center (H. Manukian et al., 2013), traveling radially outward. Interactions between the star and the SMBH can lead to exchanges in mass and momentum, which for certain geometries may result in “turbovelocity” (H. Manukian et al., 2013) or even “hypervelocity” stars (W. R. Brown, 2015; W. R. Brown et al., 2018), suggesting that many TDE remnants may be exiting the Galactic Center at elevated speeds (though see also, e.g., A. Generozov & H. B. Perets, 2022; J. J. Han et al., 2025, regarding an alternate origin). Given a selection of fast-moving stars with trajectories emanating from the Galactic Center, spectroscopic studies to look for He or N enrichment in the stellar envelope can indicate whether the candidates have experienced an extreme mixing event like a TDE. In the future, one fruitful avenue would be to simulate the expected spectra of TDE remnants to identify specific spectral fingerprints.
The introduction of the Vera C. Rubin Observatory Legacy Survey of Space and Time, coming online in the next year, provides an exciting avenue for possibly observing even more TDEs, including partial disruptions, among the expected flood of newly captured transient data. To complement the influx of new TDEs, it is important to understand what remains after the transient dims. Searching for TDE remnants requires a different strategy from capturing the bright and brief flash of a TDE itself, but we hope that studies like our own that model the detailed evolution of a TDE remnant may specify characteristics to help with the identification of TDE remnants among the stars.
@@
  • The survivor stars are initially highly altered compared to their progenitors, losing mass at a rate determined by the strength of the encounter, and increasing in radius, temperature, and luminosity, albeit for a brief period of time ($\lesssim 10^{5}$ yr). The likelihood of finding these remnants during their bright phase is low; we estimate, to within an order of magnitude, that there is roughly one such remnant in the Milky Way. During this phase, survivors have properties similar to those observed for G objects in the Galactic Center.
@@
  • Due to the intense mixing that occurs during and after the TDE, helium- and nitrogen-enriched envelopes result in the remnants. As such, spectroscopic characterization might be the most effective way of uncovering the survivor population in the Galactic Center.
@@ Appendix A Early evolution of weak encounters: A contracting envelope model
Figure 6: Left panel: Density profile of the remnant at several points in time after a $\beta=1$ encounter. During the encounter, the star becomes more diffuse and forms an extended envelope. This envelope contracts over the course of its own Kelvin–Helmholtz time. Right panel: radius and luminosity as a function of time for the first $10^{5}$ years post-disruption. The solid lines correspond to a simplified model of a contracting polytropic envelope (see text), and the dashed lines are MESA output. The rough agreement between the models supports the idea that the early “transient” evolution of the remnant corresponds to the contraction of an extended envelope that formed during the disruption as a result of energy deposition close to the surface of the star.
In Section 3 we showed that the remnants resume core hydrogen fusion on timescales $\gtrsim 10^{7}$ yr after the disruption. In this appendix, we examine the early ($\lesssim 10^{5}$ yr post-disruption) evolution of the remnant. We will discuss the $\beta=1$ encounter, although similar results are found for others.
The left panel of Figure 6 shows the density profile of the remnant at different points in time. As a result of the encounter, the envelope of the star extends to $\approx 6\,R_{\odot}$. The timescale for this extended envelope to contract noticeably is its Kelvin–Helmholtz time,
@@
The evolution of these extended envelopes has been studied in similar contexts; for example, R. Tylenda (2005) reproduced the gradual decline of the stellar merger remnant V838 Mon using a model of a contracting extended envelope. Motivated by the density structure of our TDE remnant, we will use that approach to study its early evolution. We refer the reader to R. Tylenda (2005) for a detailed construction of the model, but we summarize its main properties here.
@@
The right panel of Figure 6 compares the results of this contracting envelope model to our MESA models. The models agree reasonably well during the first ten thousand years post-disruption; the disagreement becomes significant when the radius of the envelope becomes comparable to the radius of the interior ($\approx R_{\odot}$), at which point the assumptions of the contracting envelope model begin to break down. Based on this agreement, we interpret the early evolution of the remnant as follows: The disruption deposits energy in the star. The outermost layers, which are more diffuse and loosely bound, expand significantly, then radiate the deposited energy over the course of their Kelvin–Helmholtz time. This process corresponds to the early “transient” phase of the bolometric luminosity.

### v2

@@ 4 Discussion and Conclusions
In this work, we investigate the post-disruption evolution of partial TDE remnants in order to develop strategies for discovering these objects. Our models indicate the remnant will be brightest in the first $10^{5}$ yr after disruption, during which period they may appear unusually luminous and extended compared to undisturbed stars. Over timescales of $\approx 10^{7}$$10^{8}$ yr, however, the majority of TDE remnants will have long faded to appear quite similar to main sequence stars of a similar mass. The (relatively uncommon) high-$\beta$ remnant from a more massive progenitor does appear substantially offset in luminosity and temperature from its mass-equivalent, but all remnants are still expected to fall within the natural variation of MS stars on a Hertzsprung-Russell diagram. Nevertheless, TDE remnants constitute a class of interesting objects which may be identified observationally in a few key ways.
4.1 Early Elevated Luminosity
@@
We are particularly interested in similarities between our models of TDE remnants and the known class of G objects in the Galactic Center. The G objects move with star-like kinematics, but have massive dust-shrouds that make them appear more like clouds. Although they are yet to be explained, the G objects pose interesting candidates for young TDE survivors, given their location in the Galactic Center, elevated luminosities, and puffy radii. The bolometric luminosities of G objects are commensurate with those of the brightest TDE remnants. Figure 3 compares the inferred blackbody (bolometric) luminosities of the six G objects reported in B. N. Sitarski (2016) with the early luminosity evolution of our TDE remnant models. Although their luminosity distributions are not coincident, there is some overlap between the dimmest G objects, the higher-mass TDE remnant at any age, and the lower-mass TDE remnants during their first $10^{5}$ yr post-disruption.
@@
In order to understand whether TDE remnants may be related to G objects, it is important to estimate whether TDEs occur at a rate high enough to suggest that young, high-luminosity remnants could account for some or all of the current population of G objects. Partial TDEs are expected to occur at rates up to $10\times$ higher than full disruptions; for instance, in the full loss cone regime, the rate of TDEs should scale with the inverse of $\beta$, such that low-$\beta$ encounters resulting in partial disruptions can be expected to occur much more frequently than high-$\beta$ encounters resulting in full disruptions (J. Krolik et al., 2020; N. C. Stone & B. D. Metzger, 2016; N. C. Stone et al., 2020; E. Bortolas et al., 2023). However, the heightened partial TDE occurrence rate of $\sim 10^{-4}\,\mathrm{yr}^{-1}$ is countered by the fact that our models show partial TDE remnants fading rapidly within $\lesssim 10^{5}$ yr. Accounting for both effects, we estimate that at any given moment in the Galactic Center, we can expect a handful of TDE remnants to exist that have high enough luminosities to match those of the dimmest G objects. Thus, we predict that TDE remnants in their early phase of evolution could account for a few G objects, but not the entire population (B. N. Sitarski, 2016).
To further explore the potential connection between TDE remnants and G objects, it would be worthwhile for future studies to incorporate the effect of dust. G objects have been observed to have temperatures much cooler than would be expected for their luminosities, suggesting the light of an enclosed star may be getting reprocessed to longer wavelengths by a surrounding cloud of gas and dust (e.g., G. Witzel et al., 2014). Our hydrodynamic FLASH simulation did not incorporate the effects of dust, so we cannot speak to the impact of TDEs on the production and distribution of dust in remnants, nor the dust’s subsequent evolution, but this would be a fruitful avenue for future exploration.
4.2 Kinematics & Spectroscopy
Given that the vast majority of TDE remnants have aged past their initial energized state, what are the best criteria to use when attempting to identify an older remnant after it has faded? We recommend employing a combination of kinematics and spectroscopy.
There exist several possible kinematic outcomes for TDE remnants. Disrupted stars on elliptical orbits around the SMBH will remain bound as remnants (H. Manukian et al., 2013) and may return periodically to experience repeated disruptions by the SMBH (M. MacLeod et al., 2013; M. Cufari et al., 2022; C. Liu et al., 2023; A. Bandopadhyay et al., 2024; C. Liu et al., 2025). However, those remnants that maintain parabolic or hyperbolic orbits will eventually exit the sphere of influence of the SMBH (H. Manukian et al., 2013; T. Ryu et al., 2020; M. Sharma et al., 2024), traveling radially outward. Interactions between the star and the SMBH can lead to exchanges in mass and momentum, which for certain geometries may result in “turbovelocity” stars (H. Manukian et al., 2013), suggesting that many TDE remnants may pass through the Galactic Center at speeds of a few $\times 100$ km s-1. Stars are even found in the Galactic halo moving at elevated speeds >1000>1000 km s-1, known as “hypervelocity” stars. Some hypervelocity stars may originate from the disruption of binary systems by the SMBH in the Galactic Center (J. G. Hills, 1988; H. Manukian et al., 2013; W. R. Brown, 2015; W. R. Brown et al., 2018). However, recent work analyzing the spatial and kinematic distributions of hypervelocity stars has shown that they may have an alternate origin, tracing back to the Large Magellanic Cloud instead of the Galactic Center (A. Generozov & H. B. Perets, 2022; J. J. Han et al., 2025).
What are the likely kinematics of our remnants? The stars in this study arrive on parabolic orbits according to the J. A. P. Law-Smith et al. (2020a) FLASH simulation, marginally bound to the SMBH, and receive a kick during the disruption. Previous work from H. Manukian et al. (2013) has demonstrated that the remnants of partial disruptions of single stars arriving on parabolic orbits will likely remain bound to the Milky Way, as they are not able to receive large enough kicks to result in ejection of the star from the Galaxy. H. Manukian et al. (2013) predict a population of stars in the Galactic Center with velocities that are similar to turbovelocity stars, also consistent with T. Ryu et al. (2020). From H. Manukian et al. (2013), we expect our remnants could be contributing to the population of turbovelocity stars, but likely not the hypervelocity stars, as our remnants are not unbound from the Galaxy. Our remnants will likely remain on similar orbits to their original approach, and the kick they receive due to the TDE is usually large enough that the remnants avoid a second disruption.
Thus, given a selection of fast-moving stars with trajectories emanating from the Galactic Center, we would then suggest a spectroscopic follow up study could be conducted to look for He or N enrichment in the stellar envelope. Such enrichment indicates that the candidates have experienced an extreme mixing event like a TDE. In the future, one fruitful avenue would be to simulate the expected spectra of TDE remnants to identify specific spectral fingerprints. The characteristic He and N surface enrichment of remnants will persist long after the remnants have faded from their initial luminous phase, therefore spectral signatures could be an especially useful diagnostic with the older remnant population.
4.3 Summary
@@
  • The survivor stars are initially highly altered compared to their progenitors, losing mass at a rate determined by the strength of the encounter, and increasing in radius, temperature, and luminosity, albeit for a brief period of time ($\lesssim 10^{5}$ yr). The likelihood of finding these remnants during their bright phase is low; we estimate, to within an order of magnitude, that there are roughly 1–10 such remnants in the Milky Way. During this phase, survivors have properties similar to those observed for G objects in the Galactic Center.
@@
  • Due to the intense mixing that occurs during and after the TDE, helium- and nitrogen-enriched envelopes result in all except the least-disrupted remnant. As such, spectroscopic characterization might be the most effective way of uncovering the survivor population in the Galactic Center.
The introduction of the Vera C. Rubin Observatory Legacy Survey of Space and Time, coming online in the next year, provides an exciting avenue for possibly observing even more TDEs, including partial disruptions, among the expected flood of newly captured transient data (K. Bricman & A. Gomboc, 2020). To complement the influx of new TDEs, it is important to understand what remains after the transient dims. Searching for TDE remnants requires a different strategy from capturing the bright and brief flash of a TDE itself, but we hope that studies like our own (e.g., M. Sharma et al., 2024) that model the detailed evolution of a TDE remnant may illuminate critical characteristics that will help with the identification of TDE remnants among the stars.
@@ Appendix A Early evolution of weak encounters: A contracting envelope model
Figure 6: Left panel: Density profile of the remnant at several points in time after a $\beta=1$ encounter. During the encounter, the star becomes more diffuse and forms an extended envelope. This envelope contracts over the course of its own Kelvin–Helmholtz time. Right panel: radius and luminosity as a function of time for the first $10^{5}$ yr post-disruption. The dashed lines correspond to a simplified model of a contracting polytropic envelope (see text), and the solid lines are MESA output. The rough agreement between the models supports the idea that the early “transient” evolution of the remnant corresponds to the contraction of an extended envelope that formed during the disruption as a result of energy deposition close to the surface of the star.
In this appendix, we examine the early ($\lesssim 10^{5}$ yr post-disruption) evolution of the remnant’s internal structure. In order to understand this evolution, we compare it to a simplified model of a contracting envelope. We will discuss the $1M\beta 1$ remnant, although similar results are found for the others.
The left panel of Figure 6 shows the MESA density profile of the remnant at different points in time. As a result of the encounter, the envelope of the remnant extends to $\approx 6\,R_{\odot}$. The timescale for this extended envelope to contract noticeably is its Kelvin–Helmholtz time,
@@
The evolution of these extended envelopes has been studied in similar contexts; for example, R. Tylenda (2005) reproduced the gradual decline of the stellar merger remnant V838 Mon using a model of a contracting extended envelope. Motivated by the density structure of our TDE remnant, we use that approach to study its early evolution. We refer the reader to R. Tylenda (2005) for a detailed construction of the model, but we summarize its main properties here.
@@
The right panel of Figure 6 compares this contracting envelope model to our MESA model of the remnant. The models agree reasonably well during the first ten thousand years post-disruption; the disagreement becomes significant when the radius of the envelope becomes comparable to the radius of the interior ($\approx R_{\odot}$), at which point the assumptions of the contracting envelope model begin to break down. Based on this agreement, we interpret the early evolution of the remnant as follows: The disruption deposits energy in the star. The outermost layers, which are more diffuse and loosely bound, expand significantly, then radiate the deposited energy over the course of their Kelvin–Helmholtz time. This process corresponds to the early “transient” phase of the bolometric luminosity.